OpenStudy (anonymous):

f(x)=x^4 f(x+h)-f(x)/h, f(x)-f(a)/x-a

OpenStudy (anonymous):

difference quotient problem?

OpenStudy (anonymous):

Difference Quotient: \[\frac{ f(x+h) - f(x) }{ h}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So \(f(x)=x^4\)

OpenStudy (anonymous):

do u know how to start?

OpenStudy (anonymous):

yea I do but I don't get the answer that i'm suppose to

OpenStudy (anonymous):

Oh okay, i see. Do you mind showing me your work? Or do u want me to give u an example similar your question?

OpenStudy (anonymous):

an example would be great I think im just doing something wrong

OpenStudy (anonymous):

Difference Quotient: \[\frac{ f(x+h) - f(x) }{ h }\] \[f(x) = x^2 + 2\] \[\frac{ (x+h)^2 + 2 - (x^2 + 2)}{ h }\] \[\frac{ x^2 + 2xh + h^2 + 2 - x^2 - 2 }{ h}\] \[\frac{ h(2x + h) }{ h }\] \[2x + h\] Please ask any questions you have with respect to the example.

OpenStudy (anonymous):

Everything good?

OpenStudy (anonymous):

yea i got it thanks for the help.

OpenStudy (anonymous):

Oh okay :) Glad to be of help. And good luck.